Circular No. 7675

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dk IN UGC 913 G. M. Hurst, Basingstoke, England, reports the discovery of an appar-

ent supernova (mag 15.6) by T. Boles (Coddenham, England) on an unfiltered CCD image (limiting mag ∼ 18.5) secured for the U.K. Nova/Supernova Patrol on July 29.959 UT with a 0.36-m Schmidt-Cassegrain telescope. Nothing was visible on Boles’ CCD image from 1999 Sept. 22 (limiting mag 18.5). The new object is located at α = 1^{h}22^{m}14^{s}.59, δ = +34^{o}40′11′′.3 (equinox 2000.0), which is 5′′ west and 3′′ north of the nucleus of UGC 913. M. Armstrong (Rolvenden) confirmed the presence of SN 2001dk at about the same brightness on a CCD image taken on July 31.039, which yielded the following position end figures: 14^{s}.50, 11′′.4.

SUPERNOVA 2001dl IN UGC 11725 C. Yu and W. D. Li, University of California at Berkeley, report the dis-

covery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on July 30.4 (mag ∼ 17.1) and 31.3 UT (mag ∼ 17.0). The new object is located at α = 21^{h}21^{m}01^{s}.73, δ = +9^{o}10′50′′.4 (equinox 2000.0), which is 1′′.8 west and 10′′.6 north of the nucleus of UGC 11725. A KAIT image taken on July 25.4 showed nothing at this position (limiting mag ∼ 19.0).

WZ SAGITTAE D. Steeghs and T. Marsh, Southampton University; E. Kuulkers, Space

Research Organization Netherlands and Utrecht University; and W. Skidmore, University of St. Andrews, on behalf of a larger collaboration, report: “Strong spiral waves are present in the accretion disk of WZ Sge during the first week of its outburst (IAUC 7669, 7670). Phase-resolved spectroscopy, obtained between July 28.02 and 28.24 UT with the 2.5-m Isaac Newton Telescope (+ CCD spectrograph) at La Palma, yielded 2411 spectra (resolution 5700) covering four binary orbits that were used to construct Doppler tomograms showing the distribution of line emission in the binary frame. The He II (468.6 nm) and C III (464.7 nm) images reveal that the accretion-disk emission is dominated by two spiral arms. The Hβ and He I (492.1 nm) lines, on the other hand, are dominated by highly variable absorption cores. This is the first time that spiral arms have been observed in WZ Sge, making it also the first low-mass-ratio cataclysmic variable that displays spiral arms in its disk during outburst. We encourage follow-up spectroscopy to track the evolution of these structures.”

2001 July 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green